It takes 11.2 kj of energy to raise the temperature of 145 g of benzene from 25 degrees Celsius to 70 degrees Celsius. What is the specific heat of benzene?
Please show work I am trying to figure out how to work these problems. I will rate high. Thanks
Answer #1
We know that the heat gained during raise of temperature of the substance , q = mcdt
Where m = mass of benzene = 145 g
c = spec ific heat of benzene = ?
dt = change in temoperature = 70 – 25 = 45 oC
q = heat = 11.2 KJ = 11.2 * 1000 J = 11200 J
Plug the values we get c = q / mdt
                                    = 1.716 J / g oC